### v1

# MCPVerse: An Expansive, Real-World Benchmark for Agentic Tool Use

Abstract
Large Language Models (LLMs) are evolving from text generators into reasoning agents. This transition makes their ability to use external tools a critical capability. However, evaluating this skill presents a significant challenge. Existing benchmarks are often limited by their reliance on synthetic tools and severely constrained action spaces. To address these limitations, we introduce MCPVerse, an expansive, real-world benchmark for evaluating agentic tool use. MCPVerse integrates more than 550 real-world, executable tools to create an unprecedented action space exceeding 140k tokens, and employs outcome-based evaluation with real-time ground truth for time-sensitive tasks. We benchmarked the state-of-the-art LLMs across three modes (Oracle, Standard, and Max-Scale), revealing that while most models suffer performance degradation when confronted with larger tool sets, the agentic models, such as Claude-4-Sonnet, can effectively leverage expanded exploration spaces to improve accuracy. This finding not only exposes the limitations of state-of-the-art models in complex, real-world scenarios but also establishes MCPVerse as a critical benchmark for measuring and advancing agentic tool use capabilities.
Codehttps://github.com/hailsham/mcpverse
Introduction
The ability of Large Language Models (LLMs) to interact with external tools, typically through function calling, is fundamental to their application in real-world scenarios. This capability allows them to access live data, execute code, and operate other systems, thereby moving beyond their static knowledge. Despite its importance, the evaluation of tool use is hampered by two shortcomings in current benchmarks.
First, many benchmarks rely on artificial tools (Patil et al. 2025; Li et al. 2023b). These often simulate calculators, simplified weather services, or mock shopping carts, whose data formats and interaction patterns bear little resemblance to production systems. This discrepancy allows models to succeed by recognizing superficial patterns rather than demonstrating the robust planning and coordination skills required for real-world tasks. While some benchmarks (Qin et al. 2024) claim to incorporate a wide range of real-world APIs, they often stop short of actual execution due to the complexities of live integration. Evaluation is thus confined to assessing the correctness of the selected tool name and its parameters, rather than the functional outcome of the interaction.
Second, existing benchmarks severely constrain the action space available to the model during evaluation. Even when a large set of APIs is listed, context length limitations force designers to mount only a small subset, often relying on a retrieval module to select a few dozen relevant options per query (Fei, Zheng, and Feng 2025; Li et al. 2023b; Qin et al. 2024). This strategy, while keeping prompts within the token limits of current models, prevents the assessment of a model’s ability to navigate a vast and complex solution space.
Dataset Curation MCP Support Outcome-Based Evaluation Real-Time Ground Truth Executable Tools (Total / Real / Simulated) Max Mounted Tools
MCPVerse (Ours) Human $552$ / $552$ / 0 $552$
BFCL-v3 Human $76$ / 0 / $76$ $37$
ToolBench Synthetic 0 / 0 / 0 $17$
$\tau$-bench Synthetic $28$ / 0 / $28$ $15$
API-Bank Human $73$ / 0 / $73$ 88
ToolSandBox Human $34$ / 0 / $34$ $34$
MCPBench Human $27$ / $27$ / 0 $10$
Table 1: Comparison of tool-use benchmark across key dimensions. Executable tools refers to the number of tools that may be executed during the evaluation process. Max mounted tools refers to the maximum number of tools that can be mounted for a single task.
The integration of external tools with large language models was previously hindered by a fragmented ecosystem, where different platforms relied on custom connectors and proprietary protocols. The Model Context Protocol (MCP) was introduced in 2024 as an open standard to address this challenge. By providing a uniform interface for tool access, MCP has spurred rapid adoption across major AI platforms, leading to the creation of hundreds of diverse MCP servers for applications involving web search, file systems, databases, and specialized APIs.
The widespread adoption of the MCP establishes a new paradigm for tool use. Building on this foundation, we introduce MCPVerse, a novel evaluation framework grounded in the principles of realism and scale. As shown in Table 1, MCPVerse is distinguished from prior work in three critical ways:
  • Realistic Tasks and Real-Time Verification: All tasks in MCPVerse are constructed using real-world information, such as map data and flight schedules. To handle time-sensitive queries, we developed dynamic scripts that fetch real-time ground truth, ensuring evaluation accuracy.
  • Expansive Action Space: We have carefully curated a collection of $65$ MCPs, encompassing $552$ unique tools that cover a diverse range of functionalities: file system operations, version control (Git), financial data (Yahoo Finance), news aggregation (GeekNews), lifestyle services (Amap, Variflight), office productivity (Excel), and code sandbox. The combined schemas of these tools exceed $140,000$ tokens, surpassing the context and tool-mounting limits of many state-of-the-art models and providing an unprecedentedly large action space for exploration.
  • Hybird Outcome-Based Evaluation: Recognizing that a single user request can have multiple valid solution paths, our evaluation focuses on the final outcome rather than the specific sequence of tools used. We do not penalize models for deviating from a prescribed path. We employ a hybrid evaluation method: an LLM-as-a-judge assesses the correctness of text-based outputs, while automated evaluation scripts verify state changes for tasks involving file system modifications or other environmental interactions.
Figure 1: The MCPVerse Evaluation System. A user’s task is processed by an LLM agent that can access a diverse toolset built upon MCP pool collected from MCP hubs. The agent autonomously plans and executes a sequence of function calls, and its final response is checked for correctness by an automated evaluation system.
Using our evaluation suite, we benchmarked 8 leading LLMs. Our findings show that the top-performing model, Claude-4-Sonnet, achieved an accuracy of only $57.77\%$, indicating significant room for improvement across the field. Furthermore, our experiments revealed that many state-of-the-art models are not yet equipped to handle the scale of MCPVerse, facing limitations such as a 64K context length (DeepSeek-V3), a 128-tool limit (GPT-4o-20241120), or a 512-tool limit (Gemini-2.5-Pro).
To accommodate these constraints while still probing performance at scale, we designed three evaluation modes: Oracle mode (only the pre-annotated required MCPs are provided), Standard mode (all potentially relevant MCPs are provided), and Max-Scale Mode (all 65 MCPs are mounted). Unsurprisingly, most models exhibited performance degradation as the number of available MCPs increased. More strikingly, Claude-4-Sonnet achieved a higher score in Standard mode than in the simpler Oracle mode. This counter-intuitive result underscores the significant benefit of a larger exploration space, which allows more capable models to discover more effective or robust solutions.
In summary, our contributions are:
  • We introduce MCPVerse benchmark, a comprehensive benchmark built on a large-scale set of executable, real-world tools. The benchmark features meticulously designed tasks, each accompanied by annotated ground truth or a script for time-sensitive queries.
  • We build an end-to-end, automated evaluation system that facilitates multi-turn interactions between the LLM agent and MCP tools. The system assesses the final outcome using our hybrid, outcome-based metric, which combines an LLM-as-a-judge for textual answers with dedicated scripts for verifying environmental state changes.
  • Our empirical evaluation of leading state-of-the-art models reveals their performance boundaries and practical limitations when faced with large-scale toolsets. Furthermore, our results demonstrate that a expansive action space provides a tangible benefit for agentic model, rather than merely acting as a hindrance.
The benchmark and evaluation code will be made publicly available.
Related Work
Agentic Tool Use
There is a growing interest in equipping LLMs with external tools to overcome their inherent limitations in real-world applications. Fine-tuned models such as Toolformer (Schick et al. 2023), ToolLLaMA (Qin et al. 2024), and Gorilla (Patil et al. 2024) have demonstrated the ability to accurately use different tools like calculators and search engines. In parallel, strategies such as ToolkenGPT (Hao et al. 2023), METATOOL (Wang et al. 2024b), and IPR (Xiong et al. 2024) have been introduced to enhance the compatibility and effective use of various tools. Furthermore, agentic frameworks like WebMap (Spiegel and Horák 2024), ReAct (Yao et al. 2023), Voyager (Wang et al. 2024a), Agent Reasoning (Wu, Zhu, and Liu 2025), Middleware (Gu et al. 2024), and ViperGPT (Sur´ıs, Menon, and Vondrick 2023) enable models to perform complex, multi-step tasks. These applications range from online shopping and software repair to handling multimodal inputs and learning skills in interactive environments like Minecraft.
Tool-Use Benchmarks
As LLMs are increasingly equipped with external tools to tackle complex real-world applications, the need for robust evaluation of their true capabilities has become critically important. Existing LLM benchmarks have attempted to evaluate model performance in tool-use scenarios, but as agent architectures, MCP, and diverse tool integrations have evolved, a growing gap has emerged between legacy evaluations and LLMs’ tool-use capabilities in real-world applications.
Numerous benchmarks have been developed to evaluate the tool-use capabilities of LLMs, each contributing different methodologies. Some efforts, such as ToolAlpaca (Tang et al. 2023) and NexusRaven (Srinivasan et al. 2023), have focused on automatic data generation to create large-scale evaluation sets. Another line of work has concentrated on creating diverse execution environments. For instance, Toolbench (Qin et al. 2024) provides a collection of non-executable APIs, while a more common strategy involves building sandboxed environments to simulate real-world interactions. BFCL-v3 (Patil et al. 2024), for example, implements codebases to simulate services like Twitter and mathematical calculators, and API-Bank (Li et al. 2023b) simulates common tasks such as sending emails or querying stock prices with mock data. Focusing on mobile-centric scenarios, ToolSandbox (Lu et al. 2025) simulates tools for managing device states like Wi-Fi, cellular service, and location service, while HammerBench (Wang et al. 2025) replicates API functionalities based on the features of 60 commercial applications. Other benchmarks have concentrated on specific domains, with ComplexFuncBench (Zhong et al. 2025) targeting travel-related tasks, $\tau$-bench (Yao et al. 2024) focusing on retail and airline scenarios, and ACEBench (Chen et al. 2025) covering four scenarios including food delivery and financial services. Fundamentally, these benchmarks assess model capabilities using simulated, simplified tools; their interaction logic and return values are based on fixed rules or templates, lacking the dynamism and uncertainty of real-world environments, and the complexity of the data they handle does not match that of practical applications.
While previous work has attempted to connect LLMs with real-world APIs, such as in AutoFeedback (Liu et al. 2024; Song et al. 2023), these efforts often faced challenges due to complex and non-standardized interfaces. The Model Context Protocol (MCP) (Anthropic 2024) was introduced to address this issue. MCP is a standardized protocol for LLM-tool interaction that enables dynamic discovery and orchestration of tools based on task requirements. As the protocol has matured, it has seen widespread adoption across the industry(OpenAI 2025; cur 2025; Anthropic 2025). Consequently, recent work has begun to establish benchmarks on top of MCP. For instance, MCP-Zero (Fei, Zheng, and Feng 2025) proposed a retrieval strategy for MCPs and collected an associated toolset, MCP-Tools. Meanwhile, MCPBench (Luo et al. 2025) focused on comparing the performance of different MCPs for web-search tasks.
Finally, existing benchmarks are constrained by the number of tools they can support concurrently and their restriction to non-real-time tasks. Consequently, they fail to capture the dynamic, adaptive, and complex nature of real-world agentic interactions. To address these limitations, we introduce MCP Verse, a comprehensive benchmark designed to simulate realistic, multi-step agentic workflows. MCP-Verse provides a robust evaluation framework by incorporating outcome-based evaluation, full MCP compatibility, real-time tasks, and concurrent support for up to 552 tools, thereby bridging the gap between benchmarking and real-world agent performance.
MCPVerse Design
Figure 2: Example tasks from MCPVerse. The figure shows tasks at three complexity levels, each with annotations for the required tools and the ground truth answer, which can be either static or generated by a script for time-sensitive questions.
MCP Collection
We collect a diverse set of MCPs from platforms such as Smithery, mcp.so, and ModelScope. These MCPs offer a wide range of functionalities, including local file operations, real-time flight booking, and SQL querying. As a Results, we carefully selected MCPs to be included in our benchmark based on the following criteria.
First, to ensure the long-term usability of the benchmark, we prioritize stable MCP services, such as those maintained by well-established organizations or widely adopted by the community. Second, we aim to minimize reliance on services requiring API keys. Excessive API key dependencies can introduce significant friction for researchers attempting to reproduce or extend our benchmark. Nonetheless, we retain a small number of highly valuable MCPs that do require API keys, in order to preserve the benchmark’s practical relevance. Third, we consider the evaluability of each MCP. While some MCP (e.g. Gmail, Notion) are valuable in real-world applications, they are difficult to evaluate automatically. Such MCPs are excluded from task construction but may still be included in the full MCP pool as distractors.
In total, the selected $65$ MCPs expose more than $552$ individual tools. This number may fluctuate, as MCP providers have the discretion to add or deprecate tools over time. The configuration files for all MCPs used in our benchmark are included in our codebase.
Data Curation
The collected MCP collection defines a large and diverse action space, enabling a wide variety of task types to be executed. Based on this action space, we construct a benchmark comprising $250$ tasks, which are categorized into three levels of complexity based on their required tool orchestration pattern. Figure 3 illustrates the distribution of these tasks by both complexity and time-sensitivity.
  • L1: These tasks can be completed by a single tool within $1$ or $2$ steps.
  • L2: These tasks necessitate a solution pathway of at least $5$ steps and may involve the use of a single tool or multiple tools.
  • L3: These tasks are relatively complex and may require the collaboration of different tools or the in-depth application of a specific tool, and usually take a minimum of more than 5 steps to solve.
Figure 3: Questions distribution
During the task construction phase, annotators were instructed to select interested MCPs from MCP Hubs by following the standards specified in the previous section, and then construct tasks based on these MCPs. They were required to carefully understand the definitions of each tool within the MCPs. The constructed tasks needed to meet the following characteristics: (1) There must be a solution path for the task within the selected set of MCPs; (2) The answers should be as objective and unambiguous as possible to facilitate LLM scoring; (3) The tasks should preferably be close to real-world scenarios. (4) The task should be unsolvable by the model without the use of external tools. Annotators used some MCP Client tools (e.g., Trae) to assist in question generation and answer retrieval. Each task was required to be annotated with the following attributes: question, required MCPs, required tools, time-sensitivity, complexity, and ground truth. Following the initial construction, every task underwent a comprehensive review. During this stage, we authored custom scripts to retrieve real-time answers for all time-sensitive questions.
The Choice of No Retriever
Unlike many previous tool-use benchmarks that incorporate a retrieval module (Fei, Zheng, and Feng 2025; Li et al. 2023b; Qin et al. 2024), our evaluation system deliberately omits this component for two primary reasons.
First, we hypothesize that agentic models can discover novel solution paths when presented with a large-scale action space. A retriever trained on existing patterns might inadvertently hinder this crucial exploratory potential.
Second, we anticipate a paradigm shift where LLMs, powered by the growing adoption of MCP, will function as operating systems. In this future, the number of tools connected to an LLM will vastly exceed the dozens typically seen in current benchmarks. For comparison, the average smartphone has about $80$ installed apps (buildfire 2025) (by the way, many of these apps begin to offer their own MCP services), and considering each app exposes multiple APIs (analogous to tools), the total number of available tools could easily reach the hundreds or thousands. Even with a retriever, the scale of the candidate toolset an LLM must handle will be far greater than what is benchmarked today.
Metric
Given MCPVerse’s expansive action space, there can be multiple ways to fulfill the demand, which we are not able to measure, but the end result, or end state, should be consistent with ground truth labelling. Therefore we adopt a hybrid approach for this: for tasks with textual ground truth, an LLM-as-a-judge assesses the response’s correctness, while for other tasks, such as verifying file modifications, a dedicated evaluation script is used. The scoring model we use for the LLM-as-a-judge is GPT-4o-20241120. We note that using open-source models, such as Qwen-32B, also yields a consistent ranking of results (see Appendix 2 for more details). All scoring is binary, marking the final result as either correct or incorrect.
Evaluation Results
Setup
Our evaluation system (Figure 1) is built upon the CAMEL framework (Li et al. 2023a). We aggregates all available tools from MCPs and presents them to the LLMs through a standardized function-calling mechanism. However, this unified approach introduces two significant challenges: model context length limits and navigating API-specific restrictions.
The first and most significant challenge is context length. The full MCPVerse benchmark includes $552$ tools, and their combined schemas require approximately 147k tokens. This size exceeds the context capacity of many leading models, including DeepSeek-V3 (64k), GPT-4o-20241120 (128k), and Qwen3-235B-A22B (128k). Only models with exceptionally large context windows, such as Claude-4-Sonnet (200k) and Gemini-2.5-Pro (1M), can process the entire toolset at once.
To ensure a fair and comprehensive evaluation across models, we designed three distinct modes:
  • Oracle Mode: In this setting, we load only the minimal set of MCPs required to solve a given problem. While multiple solution paths might exist, we annotate one feasible path for each question and mount only the corresponding MCPs. The context length therefore varies for each question.
  • Standard Mode: This mode is designed for the 64k-token context length, a standard supported by most state-of-the-art LLMs. We curated a set of $32$ MCPs (totaling $218$ tools), whose combined definitions occupy approximately 44k tokens. This reserves the remaining  20k tokens for the user prompt and the model’s response.
  • Max-Scale Mode: This mode loads all $65$ MCPs with $552$ tools simultaneously, which collectively occupy a context length of approximately 140k tokens. In our evaluation, only Claude-4-Sonnet and Gemini-2.5-Pro were able to complete the evaluations in this setting due to its extensive context requirements.
Model Context Tools Oracle Mode Standard Mode Max-Scale Mode
Avg. L1 L2 L3 Avg. L1 L2 L3 Avg. L1 L2 L3
DeepSeek-V3-0324 $64\mathrm{k}$ - $49.55$ $66.66$ $46.51$ $35.48$ $37.58$ $58.33$ $30.23$ $24.19$ {-} {-} {-} {-}
DeepSeek-R1-0528 $64\mathrm{k}$ - $50.77$ $61.67$ $51.94$ $38.71$ $50.28$ $58.33$ $45.74$ $46.77$ {-} {-} {-} {-}
Claude-4-Sonnet $200\mathrm{k}$ - $\mathbf{57.81}$ $68.33$ $53.49$ $\mathbf{51.61}$ $\mathbf{61.01}$ $\mathbf{70.00}$ $\mathbf{56.59}$ $\mathbf{56.45}$ $\mathbf{57.77}$ $\mathbf{66.66}$ $\mathbf{55.04}$ $\mathbf{51.61}$
Qwen3-235B-A22B $128\mathrm{k}$ - $47.90$ $63.33$ $46.51$ $33.87$ $42.26$ $58.33$ $42.64$ $25.81$ {-} {-} {-} {-}
GPT-4o-20241120 $\mathrm{k} 128128$ $51.32$ $\mathbf{72.41}$ $46.46$ $35.09$ $26.83^{*}$ $53.33^{*}$ $5.50^{*} 11$$1.67^{*}$ {-} {-} {-} {-}
Qwen3-30B-A3B $128\mathrm{k}$ - $33.31$ $53.33$ $24.03$ $22.58$ $22.78$ $36.66$ $13.95$ $17.74$ {-} {-} {-} {-}
Gemini-2.5-Pro $1\mathrm{M}$ $512$ $48.25$ $65.00$ $42.64$ $37.10$ $47.33$ $61.67$ $48.06$ $32.26$ $42.28^{*}$ $60.00^{*}$ $42.64^{*}$ $24.19^{*}$
Kimi-K2-0711 $\mathrm{k} 128128$ $55.84$ $63.33$ $\mathbf{55.81}$ $48.39$ $16.34^{*}$ $25.00^{*}$ $24.03^{*}$ $0.00^{*}$ {-} {-} {-} {-}
Table 2: Accuracy results for all evaluated models across the Oracle, Standard, and Max-Scale modes. The table also lists each model’s context length (Context) and its native tool limit (Tools). Dashes (-) indicate that a model was not evaluated in a given mode due to its context length. Scores marked with were obtained using prompt-based function calling instead of native tool use due to context length limits.
The second challenge arises from tool number limits imposed by certain APIs. For example, the native tools parameter in the GPT-4o-20241120 and Kimi-K2-0711 API is capped at $128$ tools, which is insufficient for the $218$ tools in our Standard mode. The constraint is also present at a larger scale: the Gemini-2.5-Pro API’s limit of $512$ tools is inadequate for the $552$ tools in our Max-Scale Mode. To address these limitations, we employ a method we term prompt-based function calling. This approach involves providing the complete tool definitions directly within a custom system prompt (see details in Appendix 1), rather than using the native tools parameter, thereby bypassing the API’s fixed tool count limits.
Listing 1 Evaluation Pipeline
1toolset = []
2for mcp in mcp_list:
3 mc = connect_to(mcp)
4 toolset.append(mc.get_tools())
5
6task = get_user_input()
7history = [task]
8
9while True:
10 resp = LLM(history, toolset)
11 history.append(resp)
12 if not resp.tool_calls:
13 return
14
15 for tool_call in resp.tool_calls:
16 result = call_tool(tool_call)
17 history.append(result)
Finally, we deliberately employ a simple evaluation pipeline. Our goal is to assess the fundamental capabilities of the LLM itself, rather than the effectiveness of a complex agentic framework like ReAct (Yao et al. 2023) or RolePlaying (Hu et al. 2025). These advanced methods often rely on sophisticated prompt engineering, which can obscure the model’s intrinsic abilities. Our direct approach, detailed in Listing 1, minimizes the influence of such artifacts and provides a clearer measure of the model’s core function calling proficiency. The hyperparameter settings for our experiments are detailed in the Appendix 4.
Results Analysis
This section presents a detailed analysis of the experimental results. Owing to tool-mounting constraints, the Standard mode score for GPT-4o-20241120 and the Max-Scale mode score for Gemini-2.5-Pro are both obtained via prompt-based function calling.
Overall Model Performance. Based on the results presented in Table 2, Claude-4-Sonnet demonstrates definitive leadership in the Standard mode, achieving the highest average score of $61.01$. It consistently outperforms the other models across all three complexity levels, with its advantage most pronounced on L3 tasks, where it exceeds the non–prompt-based cohort’s average by $27.10$. DeepSeek‑R1‑0528 secures a solid second place with an average score of $50.28$, closely followed by Gemini‑2.5‑Pro at $47.33$. Following this, DeepSeek-V3-0324 ($37.58$) and Qwen3-235B-A22B ($42.46$) exhibit moderate performance, positioning them in the middle tier. At the lower end of the performance spectrum are Qwen3‑30B‑A3B ($22.78$), GPT‑4o ($26.83$), and Kimi‑K2‑0711 ($16.34$). The performances of GPT‑4o and Kimi‑K2‑0711 appear to be significantly hindered by their reliance on prompt‑based function calling due to API limitations, which likely conflicts with their native tool‑use training methodologies. Meanwhile, the performance of Qwen3‑30B‑A3B is likely constrained by its smaller model size, which limits its capacity for sophisticated tool analysis.
In Oracle mode, the performance gap between models narrows. Claude-4-Sonnet maintains a lead, achieving an overall score of $57.81$, closely followed by Kimi at $55.84$. Following them, a large cohort of models—including GPT-4o-20241120, DeepSeek-R1-0528, DeepSeek-V3-0324, Gemini-2.5-Pro, and Qwen3-235B-A22B—form a tightly clustered middle tier with comparable scores, ranging from $47.90$ to $51.32$. Qwen3-30B-A3B again lags behind, a result attributable to its limited model size.
Figure 4: Model performance accuracy across different evaluation modes.
The Impact of an Expanded Action Space. As shown in Figure 4, when transitioning from the Oracle mode to the Standard mode, model performance can be categorized into four distinct groups. (1) Performance Increase: Only one model, Claude-4-Sonnet, benefits from the enlarged action space. Its average score rises from $57.77$ to $61.01$, a relative increase of $5.61\%$. The performance gains are concentrated on more difficult tasks, with improvements of $5.80\%$ on L2 and $9.38\%$ on L3, compared to a modest $2.40\%$ on L1 tasks. (2) Stable Performance: Gemini-2.5-Pro and DeepSeek-R1-0528, belong to this category, showing minimal performance changes. Their scores dropped by a negligible $0.92$ and $0.49$ points, respectively, indicating robustness to the expanded toolset. (3) Moderate Performance Drop: This group includes Qwen3-235B-A22B, Qwen3-30B-A3B, and DeepSeek-V3-0324, which exhibit a more pronounced erosion in accuracy. Their scores declined by $5.64$, $10.53$, and $11.97$ points. We attribute these declines to the increased difficulty of tool selection over a larger set and the substantially heavier context loads compared to the Oracle setting. (4) Significant Performance Drop: GPT-4o-20241120 and Kimi-K2-0711 incur the largest performance drops, with their scores falling by $24.49$ and $39.50$ points. These two models use prompt-based function calling in Standard mode, which may have a negative impact on performance and will be discussed in the following subsection.
Figure 5: A case study on the impact of an expansive action space on Claude-4-Sonnet. Blocked by a strict authentication format in the limited Oracle mode, the model succeeds in Standard mode by pivoting to an alternative tool textttfetch from the expanded toolset.
The performance increase for Claude-4-Sonnet suggests that, given sufficient context-handling capacity, the agentic model can absorb richer tool descriptions and, when blocked, actively pivot to alternative tools and strategies instead of stalling. Figure  5 presents a case study that illustrates how Claude-4-Sonnet turns an expanded toolset from a potential liability into a clear advantage. The figure shows three distinct solution pathways for the user’s task. The annotated ”correct” path (green) involves using the validate and filter tools. In the constrained Oracle mode (red path), Claude-4-Sonnet attempts this path but fails. It provides an incorrect parameter type for auth_token—passing a string instead of the expected boolean—which results in an authorization error and causes the model to abandon the task. However, in the Standard mode (blue path), Claude-4-Sonnet demonstrates adaptive problem-solving. After encountering the same initial failure, it leverages the larger available toolset to pivot to an entirely different tool, fetch. By using this alternative approach, Claude-4-Sonnet successfully resolves the problem.
The performance gap between the Oracle and Standard modes validates the effectiveness of MCPVerse. While most models perform adequately in the constrained Oracle mode, the Standard mode serves as a critical filter. It effectively distinguishes models with genuine agentic tool-use capabilities from those that falter when confronted with a larger and more complex action space.
Prompt-Based Function Calling vs. Native Function Calling. To bypass native tool number limitations, we employed a prompt-based function calling method that integrates tool schemas directly into the system prompt. A comparison (Table 3) of this prompt-based method (Prompt) with native function calling (FC) reveals two key findings.
First, Claude-4-Sonnet exhibited a significant performance degradation in both Oracle and Standard modes. We attribute this to a substantial mismatch between the function-calling templates in our prompt and those from the model’s original training. This discrepancy caused a high rate of hallucination, exceeding 70%, where the model frequently fabricated tool responses despite explicit instructions to the contrary (see Appendix 3 for details).
Model Oracle Mode Standard Mode
FC Prompt FC Prompt
Claude-4-Sonnet $57.81$ $18.33$ $\downarrow$ $61.01$ $16.65$ $\downarrow$
DeepSeek-V3-0324 $49.55$ $50.07$ $\uparrow$ $37.58$ $30.79$ $\downarrow$
Qwen3-235B-A22B $47.90$ $46.99$ $\downarrow$ $42.26$ $30.61$ $\downarrow$
Qwen3-30B-A3B $33.31$ $35.75$ $\uparrow$ $22.78$ $22.32$ $\downarrow$
GPT-4o-20241120 $51.32$ $45.95$ $\downarrow$ {-} {-}
Table 3: Performance comparison between native function calling (FC) and a prompt-based method (Prompt). The arrows indicate the performance change of the prompt-based method relative to native FC, with green signifying an improvement and red a decline.
Second, for all other models, the performance difference between the Prompt and FC methods was marginal in the constrained Oracle mode. This suggests that models can effectively execute function calls via system prompts when the number of tools is small. In the more complex Standard mode, however, these same models experienced a sharp performance decline. This indicates that as context length increases, their instruction-following capabilities weaken, and they struggle to resolve conflicts between our external formatting instructions and their internal, pre-trained templates.
Model Oracle Mode Standard Mode
L1 L2 L3 L1 L2 L3
Claude-4-Sonnet $\mathbf{7.3}$ $9.8$ $13.6$ $\mathbf{8.2}$ $\mathbf{12.3}$ $\mathbf{14.3}$
DeepSeek-V3-0324 $6.2$ $9.8$ $9.8$ $6.0$ $8.3$ $9.3$
DeepSeek-R1-0528 $6.2$ $8.3$ $11.9$ $7.2$ $9.7$ $11.8$
GPT-4o-20241120 $6.8$ $\mathbf{10.5}$ $12.2$ $5.3$ $3.8$ $4.2$
Qwen3-235B-A22B $6.0$ $8.5$ $10.7$ $5.7$ $9.0$ $8.9$
Qwen3-30B-A3B $6.3$ $8.6$ $10.8$ $6.7$ $8.7$ $7.8$
Gemini-2.5-Pro $5.2$ $6.7$ $8.3$ $5.8$ $7.1$ $7.6$
Kimi-K2-0711 $5.6$ $9.8$ $\mathbf{14.0}$ $3.8$ $4.6$ $3.2$
Table 4: Dialogue length distribution of model responses for Oracle and Standard modes.
Dialogue Length Analysis
Table 4 demonstrates the average number of dialugue length among different models across varying task complexities and tool scale. The data indicate that dialogue length scales with task complexity in both settings. In Standard mode, the mean dialogue length grows from $6.1$ (L1) to $7.9$ (L2) and $8.4$ (L3); in Oracle mode it rises from $6.2$ to $9.0$ and $11.4$, respectively.
A dialogue ends either with a solution or with abandonment after repeated failed tool calls. In Standard mode, Claude-4-Sonnet shows the longest exchanges especially on harder problems, surpassing the next-best model’s dialogue length is reduced by $2.6$ on L2 and $2.5$ on L3. These extra turns indicate a more systematic search, likely boosting its success rate, whereas other models terminate earlier on complex tasks.
Conclusion
In this work, we introduce MCPVerse, a novel benchmark designed to address two critical gaps in prior work: a lack of realism and insufficient scale. Our findings demonstrate that a large-scale action space is surprisingly advantageous for agentic models, as it empowers them to explore a comprehensive set of solution paths to find a working solution. Future work can extend the dataset’s scale and scope, as well as investigate model performance using more complex evaluation pipelines.